In physics, Ohm's law says that current through a wire, I, is directly proportional to voltage, V, and inversely proportional to resistance, R: I=V/R. It's also true that resistance is directly proportional to the length of the wire. We have a piece of wire. We pass 12 volts through this wire and measure 100 milliamps of current. If I cut the wire in half and pass 24 volts through it, how many milliamps of current will I measure?

Answer by TimothyLamb(4379)

1 milliamp = 1/1000 amp
1000 milliamp/amp
---
v = ir
r = v/i
i = v/r
---
r = 12/(100/1000)
r = 120 ohm
---
half wire resistance:
(1/2)r = 60 ohm
---
i = v/r
i = 24/60
i = 0.400 amp
i = (0.400 amp) * (1000 milliamp/amp)
i = 400 milliamp
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
Send and receive free text messages online:
https://sooeet.com